If 0.650% solution of Na2CO3 will be used to precipitate Ca+² from solution, determine the amount of solution,in grams, needed to precipitate 5.00×10² mL of 0.01120M Ca+²?


1
Expert's answer
2022-03-24T10:24:03-0400

5.00×10² mL = 5.00×10-1 L

Moles of Ca2+: 5.00 × 10-1 L * 0.01120 M = 0.0056 mol

Ca2+ + CO32- = CaCO3

The mole ratio is 1 : 1, So, 0.0056 mol of Na2CO3 is used for precipitation

Molar mass Na2CO3: 106 g/mol

Mass Na2CO3: 0.0056 * 106 = 0.5936 g

Mass of solution:

0.5936 / 0.0065 = 91.3 g